# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ VI Acknowledgments
We would like to thank Alexander Croy, Jian Han, Boris Kozinsky and Miquel Royo for helpful discussions. This work is supported by the Sinergia project of the Swiss National Science Foundation (grant number CRSII5_189924) and the NCCR MARVEL funded by the Swiss National Science Foundation (grant number 205602). S. P. acknowledges support from the Walloon Region in the strategic axe FRFS-WEL-T, the Fonds de la Recherche Scientifique de Belgique (FRS-FNRS) under grants number T.0183.23 (PDR) and T.W011.23 (PDRWEAVE). F. Macheda and F. Mauri acknowledge the MORE-TEM ERC-SYN project under grant number 951215. The computational time has been provided by the Swiss National Supercomputing Centre (CSCS) under project ID mr33, by PRACE award granting access to Discoverer in SofiaTech, Bulgaria (OptoSpin project ID 2020225411) and the IMX cluster at EPFL.
@@ Appendix B Solution of the long-wavelength equations for acoustic vibrations
$\overline{U}_{\beta}\sum_{l\kappa^{\prime}}\Phi^{l}_{\kappa\alpha,\kappa^{\prime}\beta}=0.$ (89)
@@
Since $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$ is not full-rank due to the translational invariance, there are only $33n_{\mathrm{a}}-$ 3 linearly independent first-order equations, and their solutions become also infinite and arbitrary. Without loss of generality, we remove this arbitrariness by imposing the condition of zero displacement for the first atom ($\kappa=1$) in the unit cell: $U^{(1)}_{\kappa\alpha}(\mathbf{q})=0$. The remaining $(3n_{\mathrm{a}}-3)\times(3n_{\mathrm{a}}-3)$ matrix $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$, with $\kappa$ and $\kappa^{\prime}$ both greater than one, is in general non-singular, and the solution of the first-order equation can be obtained by inverting $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$. We introduce the inverse matrix of $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$ as
@@
where $\hat{\Gamma}_{\kappa\alpha,\kappa^{\prime}\beta}$ is calculated as the pseudoinverse and satisfies the following orthogonal relationship [65]:
@@
Specifically, $\Lambda^{\kappa}_{\alpha,\beta\gamma}$ captures the force acting on the $\kappa$-th atom in the unit cell along the direction $\alpha$ due to the internal strain $\varepsilon_{\beta\gamma}$, and $\Upsilon^{\kappa}_{\alpha,\beta\gamma}$ is the corresponding atomic relaxation induced by the strain deformation.
@@
This solvability condition can be further rearranged into the equation for sound wave propagation and yields Eq. (28). Finally, the solution of the second-order equation is obtained by multiplying both sides of Eq. (99) by $\Gamma_{\kappa^{\prime\prime}\lambda,\kappa\alpha}$, summing over the indices $\kappa$ and $\alpha$, and substituting $[\omega^{(1)}(\mathbf{q})]^{2}$ with Eq. (34):
@@
which in other words implies the elastic constants $T_{\alpha\beta,\gamma\delta}$ involving the out-of-plane component are completely zero. In a nutshell, together with Eqs. (88), (98) and (107) for $U^{(0)}_{\kappa\alpha}$, $U^{(1)}_{\kappa\alpha}(\mathbf{q})$ and $U^{(2)}_{\kappa\alpha}(\mathbf{q})$, respectively, we complete the solutions of the coupled long-wavelength Eqs. (25) to (27), which are used for investigating the elastic constants of crystals in this work.
@@
with the zero-, first- and second-order solutions given by Eqs. (88), (98) and (107), respectively. To simplify the third-order solution, we can introduce the type-I in-plane strain-gradient force-response tensor due to a small deflective perturbation as
@@
The physical meaning of those high-order response tensors is that $J^{\rm{CI},\kappa}_{\alpha\beta,\gamma\delta\lambda}$ is the force acting on the atom $\kappa$ in the unit cell along the Cartesian direction $\alpha$ caused by the type-I in-plane strain gradient $\nabla_{\lambda}\tilde{\varepsilon}_{\gamma\delta}$ from the displacements of the atoms $\kappa^{\prime}$ in the out-of-plane direction $\beta$, while $J^{\rm{LM}1,\kappa}_{\alpha\gamma,\beta\delta\lambda}$ and $J^{\rm{LM}2,\kappa}_{\alpha\delta\lambda,\beta\gamma}$ represent the additional force induced by the relaxation effects of lattice under such strain-gradient perturbation. Assuming the out-of-plane direction $\beta=z$ with the origin of coordinate system placed at the mid-surface of the plate, as discussed in Sec. III and according to the Kirchhoff plate theory [106, 107, 108], the in-plane displacements vary linearly in the distance $z$ as $\mathbf{u}^{\parallel}(x,y,z)=-z\nabla w(x,y)$, from which the in-plane strain field is obtained. In addition, in order to make Eq. (109) valid and the third-order long-wavelength equation solvable, one can demonstrate that the following additional ASRs on the second-order IFCs must hold:
@@
which act as the new constraints on harmonic IFCs along with the invariance conditions in App. A. Since the left-hand side of Eq. (114) is a fifth-rank tensor, these additional ASRs contribute 60 constraint equations after considering the tensor symmetry. More generally, it is noted that the expressions involving a sublattice summation of the odd perturbation term of dynamical matrix should be always vanishing, which gives rise to such ASRs in Eqs. (86) and (114). We verify numerically in all the studied 2D materials that the violation of these extra ASRs is small and that their imposition leads to negligible changes in the values of bending rigidities. Therefore, the solution of the third-order equation can be readily obtained by inverting the matrix $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$, i.e. multiplying Eq. (69) by $\Gamma_{\kappa^{\prime\prime}\mu,\kappa\alpha}$ with a summation over $\kappa$ and $\alpha$:
@@
where $W^{\kappa}_{\alpha\beta,\gamma\delta,\lambda\mu}$ is the force-response tensor due to the Laplacian of curvature as a sum of the corresponding clamped-ion and lattice-mediated contributions, respectively. The symmetrization is performed for the two lattice-mediated tensors, $W^{\rm{LM2},\kappa}_{\alpha\beta,\gamma,\delta\lambda\mu}$ and $W^{\rm{LM3},\kappa}_{\alpha\beta,\delta\lambda\mu,\gamma}$, since they are not invariant by swapping $\gamma$ and $\delta$, while the bending rigidity tensor has such symmetry. Therefore, using the above defined sixth-rank tensors, the solvabillty condition in Eq. (116) can be recast into
@@
By multiplying $\Gamma_{\kappa^{\prime\prime}\eta,\kappa\alpha}$ on both sides of Eq. (70), summing over the indices $\kappa$ and $\alpha$, and using Eq. (75) to replace $[\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}$, the solution of the fourth-order long-wavelength equation is obtained as
@@
It is worth mentioning that Stengel [94] formulated the theory of strain-gradient elasticity using a similar variational method of lattice dynamics, where the derived strain-gradient elasticity tensor thereof for bulk solids shares many similarities with the bending rigidity tensor of this study. Our results in Eqs. (117) to (122) are fully equivalent term-by-term with Eq. (51) of Ref. [94] by summing over $\kappa$ and the symmetrization in Eq. (119). However, Stengel only considered the purely electronic and lattice-mediated terms in his real calculations, which are Eqs. (118) and (120) plus (121) with only (112) in $\Xi^{\kappa}_{\alpha\beta,\gamma\delta\lambda}$, respectively. For the bending rigidities of 2D materials, we find those remaining mixed terms (i.e. both electronic and lattice-mediated in the language of Ref. [94]) are non-negligible and crucial to recover the results of Helfrich Hamiltonian method.
@@ Appendix C Relations of the long-wavelength equations to elastic deformation
We show here how the derived long-wavelength acoustic vibrations in Eqs. (25) to (27) are closely related to the homogeneous elastic deformation of a crystal, and a similar discussion can be found in Ref. [65]. Suppose a sample whose size is smaller than the wavelength of a long acoustic wave. If one stimulates such a long wave in the sample, it will result in a homogeneous deformation with a uniform stress field and the crystal lattice will remain perfect. In this situation, the deformation can be obtained as the atomic displacements owing to the zeroth-order waves $U^{(0)}_{\alpha}(\bm{\tau})=\overline{U}_{\alpha}{\rm e}^{i\mathbf{q}\cdot\bm{\tau}}$ [65] with its derivative to position as the deformation parameter:
@@
Since we focus on a long acoustic wave ($\mathbf{q\to 0}$), the exponential can be regarded as a constant and so do $\tilde{\varepsilon}_{\alpha\beta}$, which indicates that the deformation is indeed homogeneous within the considered situation of sample. Moreover, by substituting Eqs. (22), (23) and (128) into Eq. (90), the first-order equation can be rewritten as
@@
where $U^{(1)}_{\kappa^{\prime}\beta}(\bm{\tau})=iU^{(1)}_{\kappa^{\prime}\beta}(\mathbf{q}){\rm e}^{i\mathbf{q}\cdot\bm{\tau}}$ is the first-order wave. Particularly, the right-hand side of Eq. (129) represents the force acting on the $\kappa$-th atom in unit cell due to the homogeneous strain on the crystal given by Eq. (128). Such interatomic forces resulted from the homogeneous elastic deformation create the first-order waves $U^{(1)}_{\kappa\alpha}(\bm{\tau})$, which further displace atoms to yield the internal strain that compensates the external strain fields through the left-hand side of Eq. (129), ultimately reaching a balanced strained state. Moreover, in order to further analyze the second-order waves $U^{(2)}_{\kappa\alpha}(\bm{\tau})$ which is closely related to the strain gradient field
@@
with $U^{(2)}_{\kappa^{\prime}\beta}(\bm{\tau})=U^{(2)}_{\kappa^{\prime}\beta}(\mathbf{q}){\rm e}^{i\mathbf{q}\cdot\bm{\tau}}$ defined as the second-order waves. It is easy to notice that the above equation also has the unit of forces, which corresponds to the internal force response due to the induced strain gradient field. Particularly, $\hat{T}^{\kappa}_{\alpha\beta,\gamma\delta}$ in Eq. (105) can be decomposed into the clamped-ion and lattice-mediated contributions to the internal force response due to the strain gradient and strain fields created by the zeroth-order and first-order waves, respectively. In the case of a long-wavelength deformation, the induced strain gradient field can be also assumed to be homogeneous, indicating the strain field varies linearly throughout the sample material. As can be seem from Eq. (131), the atomic displacements perturbed by such a constant strain gradient field generate the second-order waves $U^{(2)}_{\kappa\alpha}(\bm{\tau})$. When it comes to the bending waves in 2D crystals, one needs to further solve the third- and fourth-order long-wavelength equations, as shown in App. B.
@@ Appendix D Cluster expansion model for multipole and dielectric tensors
where $\bm{\kappa}=\{\kappa,i\}$ is a composite index for both the atom site $\kappa$ and the corresponding Cartesian direction $i$, $\mathcal{Q}_{\bm{\kappa}}^{I}(\alpha)$ is a flattened vector of the generalized multipole tensor of each cluster used in the expansion, $q_{I}\equiv q_{i_{1}}q_{i_{2}}\cdots q_{i_{n}}$ is a combined wavevector, and we have used the multi-index notation introduced in Ref. [163] for cluster expansion: $\alpha$ is the so-called cluster comprised of $n$ lattice sites $\{\kappa_{1},\kappa_{2},\ldots,\kappa_{n}\}$ of the primitive unit cell and $I\equiv\{i_{1},i_{2},\ldots,i_{n}\}$ is a collection of Cartesian components. For each cluster $\alpha$, its absolute value and factorial are defined as
@@
where $\alpha_{\kappa}$ is the number of the atom site $\kappa$ in the cluster. It should be noted that a cluster can contain duplicate atom sites, and such cluster is called an improper cluster, otherwise being a proper cluster. Besides, for the expansion of dielectric screening function in Eq. (133), it is actually not a cluster expansion since no atom site is involved, and we also use the introduced multi-index notation for simplicity. As can be seen from the left-hand side of Eq. (132), the cluster expansion for charge density response to a phonon perturbation only contains a single atom site $\kappa$, such an expansion is thus just an on-site cluster expansion with only the improper clusters at different orders.
@@
Table 5: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ [$e\cdot\mathrm{bohr}$] and octupole $O$ [$e\cdot\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$ and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] of BaTiO3 calculated with the same lattice structure and computational parameters as in Ref. [88], via a least-square fit to the Taylor expansion of unscreened charge density response in Eq. (43) and dielectric screening function in Eq. (II.4). In the parentheses, the components of Born effective charge and dielectric permittivity from a DFPT calculation and quadrupole, octupole and dielectric dispersion tensors from Ref. [88] are listed out for comparison. Only symmetry-independent components are shown.
@@
$\epsilon^{\infty}_{xy}$0.111 (-0.111)
$\epsilon^{(4)}_{xxxx}$29.705 (-30.159)
@@
Space-group symmetry. The charge density response $\rho_{\bm{\kappa}}(\mathbf{q})$ to a phonon perturbation is invariant under the operations in the crystal space group $\mathbb{G}_{S}$. As a result, for each symmetry operation $\hat{s}\in\mathbb{G}_{S}$, one can have the following transformation:
$\mathcal{Q}_{\bm{\kappa}}^{I}(\hat{s}\alpha)=\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha),$ (136)
where $\mathfrak{R}_{IJ}(\hat{s})=\mathfrak{r}_{i_{1}j_{1}}\otimes\cdots\otimes\mathfrak{r}_{i_{n}j_{n}}$ is a $3^{n}\times 3^{n}$ matrix which corresponds to an operation $\hat{s}$ consisting of an orthogonal transformation with a 3$\times$3 rotation matrix $\mathfrak{r}$ followed by a translation $\mathfrak{t}$, i.e. $\hat{s}\bm{\tau}_{\kappa}=\mathfrak{r}\cdot\bm{\tau}_{\kappa}+\mathfrak{t}$, and $n$ is the order of cluster $\alpha$. It is important to mention that the order of atom sites in a cluster may be changed by the symmetry operation $\hat{s}$. However, the clusters for representing multipole tensors are all improper and the order of atom sites within the cluster thus becomes trivial. Furthermore, with the help of space-group symmetry, all the clusters of lattice sites used in the expansion can be grouped into the so-called orbits which only include the symmetry-equivalent clusters of $\alpha$ under the operations of space group, denoted as $\mathbb{G}_{S\alpha}\equiv\{\hat{s}\alpha|\hat{s}\in\mathbb{G}_{S}\}$. In other words, the symmetry operations within a space group can be divided into two subgroups: one for mapping a cluster $\alpha$ to all possible clusters in its orbit $\mathbb{G}_{S\alpha}$, and the other for building symmetry constraints on the cluster $\alpha$ itself. The latter is further defined as the isotropy group $\mathbb{G}_{\alpha}$ of a cluster $\alpha$, which represents a set of transformations making $\alpha$ invariant: $\mathbb{G}_{\alpha}\equiv\{\hat{s}\in\mathbb{G}_{S}|\hat{s}\alpha=\alpha\}$. Therefore, the cluster expansion of charge density response can be recast using only the representative clusters of the symmetry-distinct atom sites as
$\overline{\rho}_{\bm{\kappa}}(\mathbf{q})=\frac{e}{\Omega}\sum_{\alpha\in\mathbb{A}/\mathbb{G}_{S}}\!\!\frac{(-i)^{|\alpha|}}{\alpha!}\!\!\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\!\!\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha)q_{I},$ (137)
where $\mathbb{A}$ stands for a set of all possible clusters, and $\mathbb{A}/\mathbb{G}_{S}$ is the orbit space that the representative clusters live in. Then, the symmetry constraints from the space group of a crystal on the multipole tensors can be built using those operations in the isotropy group of the corresponding clusters as
@@
where $\mathfrak{P}_{IJ}(\alpha)$ is a $3^{n}\times 3^{n}$ permutation matrix for a $n$-order cluster $\alpha$. One should be careful that the permutation is only allowed for the superscripted Cartesian components $I$, but not the one in the composite index $\bm{\kappa}$ which denotes the direction of phonon perturbation.

### v2

@@ VI Acknowledgments
We would like to thank Alexander Croy, Jian Han, Lev Kantorovich, Boris Kozinsky, Binbin Liu, Miquel Royo, and Yihan Wu for helpful discussions. This work is supported by the Sinergia project of the Swiss National Science Foundation (grant No. CRSII5_189924) and the NCCR MARVEL funded by the Swiss National Science Foundation (grant No. 205602). S. P. acknowledges support from the Walloon Region in the strategic axe FRFS-WEL-T, the Fonds de la Recherche Scientifique de Belgique (FRS-FNRS) under grants No. T.0183.23 (PDR) and No. T.W011.23 (PDRWEAVE). S.P. is a Research Associate of the FRS-FNRS. F. Macheda and F. Mauri acknowledge the MORE-TEM ERC-SYN project under grant No. 951215. The computational time has been provided by the Swiss National Supercomputing Centre (CSCS) under project ID mr33, by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID EHPC-EXT-2023E02-050), by PRACE award granting access to Discoverer in SofiaTech, Bulgaria (OptoSpin project ID 2020225411) and the IMX cluster at EPFL.
VII Data availability
The data and developed software that support the findings of this article are openly available [171].
@@ Appendix B Solution of the long-wavelength equations for acoustic vibrations
$\overline{U}_{\beta}\sum_{l\kappa^{\prime}}\Phi^{l}_{\kappa\alpha,\kappa^{\prime}\beta}=0,$ (89)
@@
Since $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$ is not full-rank due to the translational invariance, there are only $3n_{\mathrm{a}}-3$ linearly independent first-order equations, and their solutions also become arbitrary. Without loss of generality, we remove this arbitrariness by imposing the condition of zero displacement for the first atom ($\kappa=1$) in the unit cell: $U^{(1)}_{\kappa=1,\alpha}(\mathbf{q})=0$. The remaining $(3n_{\mathrm{a}}-3)\times(3n_{\mathrm{a}}-3)$ matrix $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$, with $\kappa$ and $\kappa^{\prime}$ both greater than one, is in general non-singular, and the solution of the first-order equation can be obtained by inverting $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$. We introduce the inverse matrix of $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$ as
@@
where $\hat{\Gamma}_{\kappa\alpha,\kappa^{\prime}\beta}$ is the inverse of the remaining $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$ matrix, and $\Gamma_{\kappa\alpha,\kappa^{\prime}\beta}$ satisfies the following orthogonal relationship [70]:
@@
[…]
@@
This solvability condition can be further rearranged into the equation for sound wave propagation and yields Eq. (28). Finally, the solution of the second-order equation is obtained by multiplying both sides of Eq. (99) by $\Gamma_{\kappa^{\prime\prime}\lambda,\kappa\alpha}$, summing over the indices $\kappa$ and $\alpha$, and substituting $\omega^{(1)}(\mathbf{q})]^{2}$ with Eq. (34):
@@
which in other words implies the elastic constants $T_{\alpha\beta,\gamma\delta}$ involving the out-of-plane component are completely zero. In a nutshell, together with Eqs. (88), (98), and (107) for $U^{(0)}_{\kappa\alpha}$, $U^{(1)}_{\kappa\alpha}(\mathbf{q})$ and $U^{(2)}_{\kappa\alpha}(\mathbf{q})$, respectively, we complete the solutions of the coupled long-wavelength Eqs. (25) to (27), which are used for investigating the elastic constants of crystals in this work.
@@
with the zero-, first-, and second-order solutions given by Eqs. (88), (98), and (107), respectively. To simplify the third-order solution, we can introduce the type-I in-plane strain-gradient force-response tensor due to a small deflective perturbation as
@@
The physical meaning of those high-order response tensors is that $J^{\rm{CI},\kappa}_{\alpha\beta,\gamma\delta\lambda}$ is the force acting on the atom $\kappa$ in the unit cell along the Cartesian direction $\alpha$ caused by the type-I in-plane strain gradient $\nabla_{\lambda}\tilde{\varepsilon}_{\gamma\delta}$ from the displacements of the atoms $\kappa^{\prime}$ in the out-of-plane direction $\beta$, while $J^{\rm{LM}1,\kappa}_{\alpha\gamma,\beta\delta\lambda}$ and $J^{\rm{LM}2,\kappa}_{\alpha\delta\lambda,\beta\gamma}$ represent the additional force induced by the relaxation effects of lattice under such strain-gradient perturbation. Assuming the out-of-plane direction $\beta=z$ with the origin of coordinate system placed at the midsurface of the plate, as discussed in Sec. III and according to the Kirchhoff plate theory [115, 116, 117], the in-plane displacements vary linearly in the distance $z$ as $\mathbf{u}^{\parallel}(x,y,z)=-z\nabla w(x,y)$, from which the in-plane strain field is obtained. In addition, in order to make Eq. (109) valid and the third-order long-wavelength equation solvable, we find that the following additional ASRs on the second-order IFCs should hold:
@@
which act as the new constraints on harmonic IFCs along with the invariance conditions in Appendix A. Since the left-hand side of Eq. (114) is a fifth-rank tensor, these additional ASRs contribute 60 constraint equations after considering the tensor symmetry. More generally, it is noted that the expressions involving a sublattice summation of the odd perturbation term of dynamical matrix should be always vanishing, which gives rise to such ASRs in Eqs. (86) and (114). We verify numerically in all the studied 2D materials that the violation of these extra ASRs is small and that their imposition leads to negligible changes in the values of bending rigidities. Therefore, the solution of the third-order equation can be readily obtained by inverting the matrix $\Phi^{(0)}_{\kappa\alpha,\kappa^{\prime}\beta}$, i.e. multiplying Eq. (69) by $\Gamma_{\kappa^{\prime\prime}\mu,\kappa\alpha}$ with a summation over $\kappa$ and $\alpha$:
@@
where $W^{\kappa}_{\alpha\beta,\gamma\delta,\lambda\mu}$ is the force-response tensor due to the Laplacian of curvature as a sum of the corresponding clamped-ion and lattice-mediated contributions, respectively. The symmetrization is performed for the two lattice-mediated tensors, $W^{\rm{LM2},\kappa}_{\alpha\beta,\gamma,\delta\lambda\mu}$ and $W^{\rm{LM3},\kappa}_{\alpha\beta,\delta\lambda\mu,\gamma}$, since they are not invariant by swapping $\gamma$ and $\delta$, while the bending rigidity tensor has such symmetry. Therefore, using the above defined sixth-rank tensors, the solvability condition in Eq. (116) can be recast into
@@
By multiplying $\Gamma_{\kappa^{\prime\prime}\eta,\kappa\alpha}$ on both sides of Eq. (70), summing over the indices $\kappa$ and $\alpha$, and using Eq. (75) to replace $\omega^{(2)}_{\rm ZA}(\mathbf{q})]^{2}$, the solution of the fourth-order long-wavelength equation is obtained as
@@
It is worth mentioning that Stengel [101] formulated the theory of strain-gradient elasticity using a similar variational method of lattice dynamics, where the derived strain-gradient elasticity tensor thereof for bulk solids shares many similarities with the bending rigidity tensor of this study. Our results in Eqs. (117) to (122) are fully equivalent term-by-term with Eq. (51) of Ref. [101] by summing over $\kappa$ and the symmetrization in Eq. (119). However, Stengel only considered the purely electronic and lattice-mediated terms in his real calculations, which are Eqs. (118) and (120) plus (121) with only (112) in $\Xi^{\kappa}_{\alpha\beta,\gamma\delta\lambda}$, respectively. For the bending rigidities of 2D materials, we find those remaining mixed terms (i.e. both electronic and lattice-mediated in the language of Ref. [101]) are non-negligible and crucial to recover the results of the Helfrich Hamiltonian method.
@@ Appendix C Relations of the long-wavelength equations to elastic deformation
We show here how the derived long-wavelength acoustic vibrations in Eqs. (25) to (27) are closely related to the locally homogeneous elastic deformation of a crystal, and a similar discussion can be found in Ref. [70]. Suppose a sample whose size is smaller than the wavelength of a long acoustic wave. If one stimulates such a long wave in the sample, it will result in a homogeneous deformation with a uniform stress field and the crystal lattice will remain perfect. In this situation, the deformation can be obtained as the atomic displacements owing to the zeroth-order waves $U^{(0)}_{\alpha}(\bm{\tau})=\overline{U}_{\alpha}{\rm e}^{i\mathbf{q}\cdot\bm{\tau}}$ [70] with its derivative to position as the deformation parameter:
@@
Since we focus on a long acoustic wave ($\mathbf{q\to 0}$), the exponential can be regarded as a constant and so do $\tilde{\varepsilon}_{\alpha\beta}$, which indicates that the deformation is indeed homogeneous within the considered situation of sample. Moreover, by substituting Eqs. (22), (23), and (128) into Eq. (90), the first-order equation can be rewritten as
@@
[…]
@@
with $U^{(2)}_{\kappa^{\prime}\beta}(\bm{\tau})=U^{(2)}_{\kappa^{\prime}\beta}(\mathbf{q}){\rm e}^{i\mathbf{q}\cdot\bm{\tau}}$ defined as the second-order waves. It is easy to notice that the above equation also has the unit of forces, which corresponds to the internal force response due to the induced strain gradient field. Particularly, $\hat{T}^{\kappa}_{\alpha\beta,\gamma\delta}$ in Eq. (105) can be decomposed into the clamped-ion and lattice-mediated contributions to the internal force response due to the strain gradient and strain fields created by the zeroth-order and first-order waves, respectively. In the case of a long-wavelength deformation, the induced strain gradient field can be also assumed to be homogeneous, indicating the strain field varies linearly throughout the sample material. As can be seem from Eq. (131), the atomic displacements perturbed by such a constant strain gradient field generate the second-order waves $U^{(2)}_{\kappa\alpha}(\bm{\tau})$. When it comes to the bending waves in 2D crystals, one needs to further solve the third- and fourth-order long-wavelength equations, as shown in Appendix B.
@@ Appendix D Cluster expansion model for multipole and dielectric tensors
where $\bm{\kappa}\equiv\{\kappa,i\}$ is a composite index for both the atom site $\kappa$ and the corresponding Cartesian direction $i$, $\mathcal{Q}_{\bm{\kappa}}^{I}(\alpha)$ is a flattened vector of the generalized multipole tensor of each cluster used in the expansion, $q_{I}\equiv q_{i_{1}}q_{i_{2}}\cdots q_{i_{n}}$ is a combined wave vector, and we have used the multi-index notation introduced in Ref. [172] for cluster expansion: $\alpha$ is the so-called cluster comprised of $n$ lattice sites $\{\kappa_{1},\kappa_{2},\ldots,\kappa_{n}\}$ of the primitive unit cell and $I\equiv\{i_{1},i_{2},\ldots,i_{n}\}$ is a collection of Cartesian components. The order of a cluster is also the number of atom sites within it. For each cluster $\alpha$, its absolute value and factorial are defined as
@@
where $\alpha_{\kappa}$ is the number of the atom site $\kappa$ in the cluster. It should be noted that a cluster can contain duplicate atom sites, and such a cluster is called an improper cluster, otherwise being a proper cluster. As can be seen from the left-hand side of Eq. (132), the cluster expansion for charge density response to a phonon perturbation only contains a single atom site $\kappa$, and such an expansion is thus just an on-site cluster expansion with only the improper clusters at different orders, i.e. the cluster $\alpha$ comprising only the repeated atom sites $\kappa$ in the charge density response. For a given atom site $\kappa$, we should also stress that the cluster $\alpha$ only defines the superscript components II of a multipole tensor, which actually represent the electric field response, and the order of multipole tensor is hence $n+1$ given an $n$-order cluster. Besides, for the expansion of dielectric screening function in Eq. (133), it is actually not a cluster expansion since no atom site is involved, and we also use the introduced multi-index notation for simplicity.
@@
Table 5: The dynamical Born effective charge $Z$ [$e$], quadrupole $Q$ $e~\mathrm{bohr}$] and octupole $O$ $e~\mathrm{bohr}^{2}$], clamped-ion dielectric permittivity $\bm{\epsilon}^{\infty}$ and dielectric dispersion tensors $\bm{\epsilon}^{(4)}$ [bohr2] of BaTiO3 calculated with the same lattice structure and computational parameters as in Ref. [95], via a least-squares fit to the Taylor expansion of unscreened charge density response in Eq. (43) and dielectric screening function in Eq. (II.4). In the parentheses, the components of Born effective charge and dielectric permittivity from a DFPT calculation and quadrupole, octupole and dielectric dispersion tensors from Ref. [95] are listed out for comparison. Only symmetry-independent components are shown.
@@
$\epsilon^{\infty}_{xy}$ -0.111 (-0.111)
$\epsilon^{(4)}_{xxxx} -$29.705 (-30.159)
@@
Space-group symmetry. The charge density response $\overline{\rho}_{\bm{\kappa}}(\mathbf{q})$ to a phonon perturbation transforms covariantly in its atom site $\kappa$ under the operations in the crystal space group $\mathbb{G}_{S}$. As a result, for each symmetry operation $\hat{s}\in\mathbb{G}_{S}$, one can have the following transformation:
$\mathcal{Q}_{\bm{\kappa}^{\prime}}^{I}(\hat{s}\alpha)=\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha),$ (136)
where $\mathfrak{R}_{IJ}(\hat{s})=\mathfrak{r}_{i_{1}j_{1}}\cdots\mathfrak{r}_{i_{n+1}j_{n+1}}$ is a $3^{n+1}\times 3^{n+1}$ matrix, which corresponds to an operation $\hat{s}$ consisting of an orthogonal transformation with a $\times3$ rotation matrix $\mathfrak{r}$ followed by a translation $\mathfrak{t}$, i.e. $\hat{s}\bm{\tau}_{\kappa}=\mathfrak{r}\cdot\bm{\tau}_{\kappa}+\mathfrak{t}$, and $n$ is the order of cluster $\alpha$. It is important to mention that the order of atom sites in a cluster may be changed by the symmetry operation $\hat{s}$. However, the clusters for representing multipole tensors are all improper and the order of atom sites within the cluster thus becomes trivial. Furthermore, with the help of space-group symmetry, all the clusters of lattice sites used in the expansion can be grouped into the so-called orbits which only include the symmetry-equivalent clusters of $\alpha$ under the operations of space group, denoted as $\mathbb{G}_{S\alpha}\equiv\{\hat{s}\alpha|\hat{s}\in\mathbb{G}_{S}\}$. In other words, the symmetry operations within a space group can be divided into two subgroups: one for mapping a cluster $\alpha$ to all possible clusters in its orbit $\mathbb{G}_{S\alpha}$, and the other for building symmetry constraints on the cluster $\alpha$ itself. The latter is further defined as the isotropy group $\mathbb{G}_{\alpha}$ of a cluster $\alpha$, which represents a set of transformations making $\alpha$ invariant: $\mathbb{G}_{\alpha}\equiv\{\hat{s}\in\mathbb{G}_{S}|\hat{s}\alpha=\alpha\}$. Therefore, the cluster expansion of charge density response can be recast using only the representative clusters of the symmetry-distinct atom sites as
$\overline{\rho}_{\bm{\kappa}^{\prime}}(\mathbf{q})=\frac{e}{\Omega}\frac{(-i)^{|\alpha|}}{\alpha!}\sum_{\hat{s}\alpha\in\mathbb{G}_{S\alpha}}\mathfrak{R}_{IJ}(\hat{s})\mathcal{Q}_{\bm{\kappa}}^{J}(\alpha)q_{I}.$ (137)
Then, the symmetry constraints from the space group of a crystal on the multipole tensors can be built using those operations in the isotropy group of the corresponding clusters as
@@
where $\mathfrak{P}_{IJ}(\alpha)$ is a $3^{n+1}\times 3^{n+1}$ permutation matrix for a $n$-order cluster $\alpha$. One should be careful that the permutation is only allowed for the superscripted Cartesian components $I$, but not the one in the composite index $\bm{\kappa}$ which denotes the direction of phonon perturbation.